40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 31 23:12:17 2018, gautam, Update, LSC, DRMI locking attempt REFL55_whtCheck.pdf
    Reply  Fri Jun 1 01:22:50 2018, gautam, Update, LSC, DRMI locking restored DRMI_20180531.png
       Reply  Fri Jun 8 01:58:56 2018, gautam, Update, LSC, DRMI locking attempt again 
          Reply  Tue Jun 12 03:22:25 2018, gautam, Update, LSC, AUX laser shuttered 
             Reply  Wed Jun 13 01:02:40 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking 
                Reply  Wed Jun 13 11:17:40 2018, gautam, Update, LSC, PRCL loop shape anomaly PRCL_12Jun2018_WeirdShape.pdf
                   Reply  Thu Jun 14 00:40:42 2018, gautam, Update, LSC, PRCL loop shape anomaly 
                Reply  Thu Jun 14 18:09:24 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking DRMI1f_June14.pdf
                   Reply  Fri Jun 15 00:53:21 2018, gautam, Update, LSC, Calibrated MICH spectrum C1NB_disp_40m_MICH_NB_2018-06-14.pdf
Message ID: 13907     Entry time: Thu May 31 23:12:17 2018     Reply to this: 13908
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking attempt 

Summary:

I wanted to recover the DRMI locking. Among other things, Jon mentioned that his mode spectroscopy can be done in the DRMI config. But I was foiled last night by a rogue waveplate in the AS beampath, and today evening, I noticed the resurfacing of this problem. Clearly, this is indicative of some issue in the analog whitening electronics, as the DC light level on the AS55 PD is consistent with previous measurements. Moreover, last time, the problem "fixed itself" so I don't know what exactly the problem was in the first place. I'll try doing the same test in the linked elog tomorrow. As a quick test, I cycled through the whitening gains (0-45dB) to see if it was some stuck ADC register, but that didn't fix the problem.

The problem seems to be with REFL55 only - I am able to lock the PRMI with carrier resonant without any issues, and the error signal levels are consistent with what I remember them being while the PRMI is swinging around. AS55 lives on the same whitening board and doesn't seem to suffer from the same probelms.


Decided to do the check tonight, but as Attachment #1 shows, no real red flags from the whitening gain side.

Attachment 1: REFL55_whtCheck.pdf  27 kB  Uploaded Fri Jun 1 01:16:27 2018  | Hide | Hide all
REFL55_whtCheck.pdf
ELOG V3.1.3-